# On obstructions to the Euler system method for Rankin-Selberg convolutions

### v2

@@ 2 Special elements in the image of the convolution of two modular forms
The matrix $N$ has finite order, so it is similar to a diagonal matrix $\Delta_{u,v}$ where $u,v$ are roots of unity. In this case, after renaming if needed, it is necessary that $u\neq v$, and so $\ker(M-u^{-1}I_{2})$ is of rank one. Since $\det{M}$ is a fixed root of unity $\delta$, the two eigenvalues of $M$ are $u^{-1}$ and $\delta u$. The question then becomes: can we find a non-scalar $M\in\alpha\operatorname{GL}_{2}(\mathcal{O}_{F_{\mathfrak{p}}})$ such that its eigenvalues are $u^{-1}$ and $\delta u$? There is at least one case where we can answer in the affirmative: assuming that $u,\delta,\alpha\in F_{\mathfrak{p}}$, we can choose $M=\Delta_{u^{-1},\delta u}$ unless $u^{-1}=\delta u$, in which case we can take $M=u^{-1}U$.
@@
If $u^{2}\neq\varepsilon_{f}(\sigma)^{-1}$, by Corollary 2, we can choose $\sigma^{\prime}\in\sigma G_{H(\mu_{p^{\infty}})}$ such that it acts on $T_{f,\mathfrak{p}}$ as $\Delta_{u^{-1},u\varepsilon_{f}(\sigma)}$ and on $T_{g}$ by $\rho_{g}(\sigma)$. Then the eigenvalues of $\sigma^{\prime}$ on $T_{f,g,\mathfrak{p}}$ are $1,u^{2}\varepsilon_{f}(\sigma),(\varepsilon_{f}\varepsilon_{g})(\sigma),v/u$: $1$ is distinct from all the others, hence (by our choice of $p$), not congruent to any of the others mod $\mathfrak{p}$, and $\sigma^{\prime}$ works.
@@
The above Proposition is not sufficient in order to answer Question Introduction: for general choices of $\alpha,u$, we may not assume that they are contained in $F_{\mathfrak{p}}$. There is however one special case in which the argument still applies: the case where $\alpha=1$ and $u=\pm 1$. The goal of the two following Propositions is to give sufficient conditions that imply that $\rho_{g}(G_{H})$ contains a non-scalar matrix with ±$\pm 1$ as an eigenvalue.
@@
special2special Proposition 2.6  Assume that $-1\in\varepsilon_{g}(G_{H})$ and that $\rho_{g}(G_{H})$ does not contain an element similar to $\Delta_{-1,1}$. Then $g$ has either real or complex multiplication by some quadratic field $K$. If moreover $K\not\subset H$ for some choice of $K$, then the following holds:
@@
Proof. –  Let $M\in\operatorname{GL}_{2}(L_{0})$ be a matrix with $\det{M}=-1$ whose image in $\operatorname{PGL}_{2}(L_{0})$ has order $2$. Then $M^{2}$ is scalar, so $M$ is conjugate in $\operatorname{GL}_{2}({\mathbb{C}})$ to $\Delta_{a,b}$ for some $a,b\in L_{0}^{\times}$ with $a=-b$ and $ab=-1$. Therefore $M$ is similar to $\Delta_{1,-1}$ in $\operatorname{GL}_{2}({\mathbb{C}})$, so $M$ is similar to $\Delta_{1,-1}$ in $\operatorname{GL}_{2}(L_{0})$.
Let $Z\leq\rho_{g}(G_{H})$ be the subgroup of scalar matrices. Consider the surjective determinant map $\delta:\tilde{\rho}_{g}(G_{H})=\rho_{g}(G_{H})/Z\rightarrow\varepsilon_{g}(G_{H})/\det{Z}$, which has cyclic image. Let $x\in\tilde{\rho}_{g}(G_{H})$ be an element of order $2$ such that $\delta(x)=-\det{Z}$: then there is a matrix $M\in\rho_{g}(G_{H})$ above $x$ such that $\det{M}=-1$, so $M$ is similar to $\Delta_{1,-1}$, whence a contradiction.
@@
Case 1: when $-1\notin\det{Z}$.
@@
The image of $\delta$ contains exactly two elements of $2$-torsion: they are $-\det{Z}$ and the trivial element $\det{Z}$. For any $x\in\tilde{\rho}_{g}(G_{H})$ of order two, $\delta(x)\in\mathrm{im}(\delta)$ is a $2$-torsion element which is not $-\det{Z}$, so $x\in\ker{\delta}$. Since $\tilde{\rho}_{g}(G_{H})/\ker{\delta}$ contains an element of order two, the subgroup of $\tilde{\rho}_{g}(G_{H})$ generated by its elements of order two has even index: thus, $\tilde{\rho}_{g}(G_{H})$ is not dihedral, and it is not isomorphic to $A_{4}$, $S_{4}$ or $A_{5}$ either. Since $\tilde{\rho}_{g}(G_{H})$ is a subgroup of $\operatorname{PGL}_{2}(L_{0})$, it is cyclic with order $n$ divisible by $4$.
@@
Therefore, the subgroup $\tilde{\rho}_{g}(G_{{\mathbb{Q}}})^{\prime}$ of $\tilde{\rho}_{g}(G_{H})$ is cyclic, so $\tilde{\rho}_{g}(G_{{\mathbb{Q}}})$ is not isomorphic to $A_{4}$, $S_{4}$, or $A_{5}$. Hence $\tilde{\rho}_{g}(G_{H})$ is dihedral and $g$ has either real or complex multiplication by a quadratic number field $K$. The dihedral group $\tilde{\rho}_{g}(G_{{\mathbb{Q}}})$ contains the cyclic subgroup $\tilde{\rho}_{g}(G_{H})$ of order $n\geq 4$, so $\tilde{\rho}_{g}(G_{H})$ is contained in the unique index two cyclic subgroup of $\tilde{\rho}_{g}(G_{{\mathbb{Q}}})$ (whose inverse image by $\tilde{\rho}_{g}$ is $G_{K}$), so that $K$ is unique and one has $K\subset H$, so we are done.
Case 2: when $-1\in\det{Z}$.
@@
In this case, $\rho_{g}(G_{H})$ contains the scalar matrix $iI_{2}$. Then the group $\tilde{\rho}_{g}(G_{{\mathbb{Q}}})^{\prime}\subset\ker{\delta}$ does not contain an element of order $2$, so it has odd cardinality. This is not possible if the group $\tilde{\rho}_{g}(G_{{\mathbb{Q}}})$ is isomorphic to $A_{4}$, $S_{4}$ or $A_{5}$ or if it is dihedral with cardinality divisible by $8$. Hence $\tilde{\rho}_{g}(G_{{\mathbb{Q}}})$ is dihedral with cardinality not divisible by $8$, so in particular, $g$ has either real or complex multiplication by some quadratic number field $K$.
@@
Let us furthermore assume that $K\not\subset H$. Let $C$ denote the cyclic subgroup of $D:=\tilde{\rho}_{g}(G_{{\mathbb{Q}}})$ such that $G_{K}=\tilde{\rho}^{-1}(C)$, and $D_{H}$ denote the subgroup $\tilde{\rho}_{g}(G_{H})$ of $D$. Since $K\not\subset H$, one has $D_{H}\not\subset C$, so $D_{H}$ is dihedral with cyclic subgroup $D_{H}\cap C$ of index $2$ (it is possible, given our assumptions, that $D_{H}\simeq{\mathbb{Z}}/2{\mathbb{Z}}$ and $D_{H}\cap C$ is trivial; this case does not require any special treatment). In particular, the abelianization of $D_{H}$ has exponent $2$, so the same holds for its cyclic quotient $\mathrm{im}(\delta)$, thus $\mathrm{im}(\delta)$ has at most two elements, i.e. either one has $\det{Z}=\varepsilon_{g}(G_{H})$ or one has $\det{Z}=\varepsilon_{g}(G_{H})^{2}$.
@@
Let $x\in D_{H}\backslash C$: then $x$ has order $2$, so $\delta(x)\neq-\det{Z}=\det{Z}$, so the image of $\delta$ has two elements, hence $-1\in\det{Z}=\varepsilon_{g}(G_{H})^{2}$, i.e. $i\in\varepsilon_{g}(G_{H})$. Furthermore, $\ker{\delta}$ is a subgroup of index $2$ of $D_{H}$ without an element of order two, so it has odd cardinality, whence the cardinality of $\tilde{\rho}_{g}(G_{H})$ is not divisible by $4$Box
@@
    $q=3$ divides $[H:{\mathbb{Q}}]$, $\tilde{\rho}_{g}(G_{{\mathbb{Q}}})$ is isomorphic to $A_{4}$, $\tilde{\rho}_{g}(G_{H})$ is isomorphic to ${\mathbb{F}}_{2}^{\oplus 2}$.
@@
Consider the determinant map $\delta:\rho_{g}(G_{H})\rightarrow\varepsilon_{g}(G_{H})/\det{Z}$. It is surjective; its kernel contains $Z$ and $\rho_{g}(G_{{\mathbb{Q}}})^{\prime}$, since the latter is contained in $\operatorname{SL}_{2}(L_{0})$: hence $\delta$ factors through the quotient $\rho_{g}(G_{H})/Z\rho_{g}(G_{{\mathbb{Q}}})^{\prime}\simeq\tilde{\rho}_{g}(G_{H})/\tilde{\rho}_{g}(G_{{\mathbb{Q}}})^{\prime}$; in other words, $\varepsilon_{g}(G_{H})/\det{Z}$ identifies as a cyclic quotient of $\tilde{\rho}_{g}(G_{H})/\tilde{\rho}_{g}(G_{{\mathbb{Q}}})^{\prime}$.
Case 1: $[\tilde{\rho}_{g}(G_{{\mathbb{Q}}}):\tilde{\rho}_{g}(G_{{\mathbb{Q}}})^{\prime}]$ is prime to $q$.
In this case, every element of $\tilde{\rho}_{g}(G_{{\mathbb{Q}}})$ with order $q$ is contained in $\tilde{\rho}_{g}(G_{{\mathbb{Q}}})^{\prime}$. Because $H/{\mathbb{Q}}$ is abelian, the group $\rho_{g}(G_{{\mathbb{Q}}})^{\prime}=\rho_{g}(G_{{\mathbb{Q}}}^{\prime})$ is contained in $\rho_{g}(G_{H})$ and in $\operatorname{SL}_{2}(L_{0})$, and its image in $\operatorname{PGL}_{2}(L_{0})$ is equal to $\tilde{\rho}_{g}(G_{{\mathbb{Q}}})^{\prime}$. Hence the image of $\rho_{g}(G_{H})\cap\operatorname{SL}_{2}(L_{0})$ in $\operatorname{PGL}_{2}(L_{0})$ contains an element of order $q$.
@@
Case 2: $[\tilde{\rho}_{g}(G_{{\mathbb{Q}}}):\tilde{\rho}_{g}(G_{{\mathbb{Q}}})^{\prime}]$ is divisible by $q$.
If $\tilde{\rho}_{g}(G_{{\mathbb{Q}}})$ is isomorphic to $S_{4},A_{5}$ or is dihedral, then $\tilde{\rho}_{g}(G_{{\mathbb{Q}}})/\tilde{\rho}_{g}(G_{{\mathbb{Q}}})^{\prime}$ is a $2$-group, so its cardinality is not divisible by $q$. Hence $\tilde{\rho}_{g}(G_{{\mathbb{Q}}})$ is isomorphic to $A_{4}$; since $\tilde{\rho}_{g}(G_{{\mathbb{Q}}})/\tilde{\rho}_{g}(G_{{\mathbb{Q}}})^{\prime}$ is isomorphic to ${\mathbb{Z}}/3{\mathbb{Z}}$, one also has $q=3$.
@@
Let $M\in\rho_{g}(G_{H})$ be such that $\det{M}$ is a primitive third root of unity. If $M^{2}\in Z$, $M^{4}=\det(M^{2})I_{2}\in Z$ is an element of order $3$, a contradiction: hence the projective image of $M$ has order at least $3$. The elements of $A_{4}$ have order $1$, $2$ or $3$, so the projective image of $M$ has order $3$, i.e. $M^{3}\in\operatorname{SL}_{2}(L_{0})$ is scalar, thus $M^{3}\in\{\pm I_{2}\}$. Replacing $M$ with $M^{2}$ if needed, we may assume that $M^{3}=I_{2}$. Let $u,v\in L_{0}$ be the two eigenvalues of $M$, then one has $u^{3}=v^{3}=1$ and $u\notin\{v,v^{-1}\}$ (since $M$ is not scalar and $\det{M}$ is a primitive third root of unity), so exactly one of $u$ and $v$ is equal to $1$, and we are done.
@@
If $Z$ does not contain any element of order $3$, we get back to Case 2 (b). So we may assume that $Z$ contains an element of order $3$. This implies that the cardinalities of the cyclic groups $\varepsilon_{g}(G_{H})$ and $\det{Z}$ are divisible by $3$ and not by $9$, so the group $\varepsilon_{g}(G_{H})/\det{Z}$ has cardinality $d$ coprime to $3$. Let $M\in\rho_{g}(G_{H})$ be such that its image in $\tilde{\rho}_{g}(G_{H})\simeq A_{4}$ has order $3$. Then $(\det{M})^{d}\in\det{Z}$, so there exists $\lambda\in L_{0}^{\times}$ such that $\lambda I_{2}\in\rho_{g}(G_{H})$ and $(\det{M})^{d}=\lambda^{2}$. Hence, the projective image of $M_{1}=(\lambda I_{2})^{-1}M^{d}\in\rho_{g}(G_{H})\cap\operatorname{SL}_{2}(L_{0})$ has order $3$. Since $Z$ contains elements of order $3$, we are done by the argument at the beginning of the proof.
@@
sufficient-for-euler-adaptedspecial Corollary 2.8  Let $\mathfrak{p}$ be a good prime for $(f,g)$. If $\rho_{g}(G_{H})$ contains a non-scalar matrix $M$ such that $1$ or $-1$ is an eigenvalue of $M$, then $T_{f,g,\mathfrak{p}}$ satisfies (wE); if furthermore the order of $M$ is prime to $p$, then $T_{f,g,\mathfrak{p}}$ satisfies (sE). Hence $T_{f,g,\mathfrak{p}}$ has Euler-adapted Galois image for all but finitely many primes $\mathfrak{p}$. In particular, all of the above applies if any of the following conditions is satisfied:
@@
    sc2special $-1\in\varepsilon_{g}(G_{H})$ and $\tilde{\rho}_{g}(G_{{\mathbb{Q}}})$ is not dihedral.
@@
    sc3special $-1\in\varepsilon_{g}(G_{H})$, $\tilde{\rho}_{g}(G_{H})$ is a dihedral group of order $4n$ with $n\geq 1$.
@@
    sc4special $-1\in\varepsilon_{g}(G_{H})$, $\tilde{\rho}_{g}(G_{{\mathbb{Q}}})$ is a dihedral group of order $8n$ with $n\geq 1$, and $g$ has either real or complex multiplication by some quadratic field $K\not\subset H$.
@@
    sc5special $\varepsilon_{g}(G_{H})$ contains $-1$ but not $i$, and, for some quadratic number field $K$ by which $g$ has either real or complex multiplication, one has $K\not\subset H$.
@@
Proof. –  By Proposition 2, we only need to discuss conditions (wE) and (sE). Let $\mathfrak{p}$ be a good prime with residue characteristic $p$ and $M\in\rho_{g}(G_{H})$ be a non-scalar matrix admitting $s\in\{\pm 1\}$ as an eigenvalue. By Proposition 2, there exists $\sigma\in G_{H(\mu_{p^{\infty}})}$ such that $\rho_{g}(\sigma)=M$ and $\sigma$ acts on $T_{f,\mathfrak{p}}$ by $sU$. Then the kernel of the action of $\sigma-1$ on $T_{f,g,\mathfrak{p}}\otimes(L_{0})_{\mathfrak{p}}$ is a line and $T_{f,g,\mathfrak{p}}$ satisfies (sE). If the order of $M$ is prime to $p$, then $\mathcal{O}_{(L_{0})_{\mathfrak{p}}}^{\oplus 2}/(M-sI_{2})\mathcal{O}_{(L_{0})_{\mathfrak{p}}}^{\oplus 2}$ is a free $\mathcal{O}_{(L_{0})_{\mathfrak{p}}}$-module of rank one, whence $T_{f,g,\mathfrak{p}}/(\sigma-1)T_{f,g,\mathfrak{p}}$ is a free $\mathcal{O}_{(L_{0})_{\mathfrak{p}}}$-module of rank one, thus $T_{f,g,\mathfrak{p}}$ satisfies (sE).
All that remains to do is show that in conditions (i)-(vi), there exists a non-scalar matrix $M\in\rho_{g}(G_{H})$ of which $1$ or $-1$ is an eigenvalue.

### v3

@@ 2 Special elements in the image of the convolution of two modular forms
The matrix $N$ has finite order, so it is similar to a diagonal matrix $\Delta_{u,v}$ where $u,v$ are roots of unity. In this case, after renaming if needed, it is necessary that $u\neq v$ and that $\ker(M-u^{-1}I_{2})$ has rank one. Since $\det{M}$ is a fixed root of unity $\delta$, the two eigenvalues of $M$ are $u^{-1}$ and $\delta u$. The question then becomes: can we find a non-scalar $M\in\alpha\operatorname{GL}_{2}(\mathcal{O}_{F_{\mathfrak{p}}})$ such that its eigenvalues are $u^{-1}$ and $\delta u$? There is at least one case where we can answer in the affirmative: assuming that $u,\delta,\alpha\in F_{\mathfrak{p}}$, we can choose $M=\Delta_{u^{-1},\delta u}$ unless $u^{-1}=\delta u$, in which case we can take $M=u^{-1}U$.
@@
If $u^{2}\neq\varepsilon_{f}(\sigma)^{-1}$, by Corollary 2, we can choose $\sigma^{\prime}\in\sigma G_{H(\mu_{p^{\infty}})}$ such that it acts on $T_{f,\mathfrak{p}}$ as $\Delta_{u^{-1},u\varepsilon_{f}(\sigma)}$ and on $T_{g}$ by $\rho_{g}(\sigma)$. Then the eigenvalues of $\sigma^{\prime}$ on $T_{f,g,\mathfrak{p}}$ are $u^{2}\varepsilon_{f}(\sigma),(\varepsilon_{f}\varepsilon_{g})(\sigma),v/u: 11$ is distinct from all the others, hence (by our choice of $p$), not congruent to any of the others mod $\mathfrak{p}$, and $\sigma^{\prime}$ works.
@@
The above Proposition is not sufficient in order to answer Question Introduction: in general, we cannot assume that we can find $\alpha,u$ that are both contained in $F_{\mathfrak{p}}$. There is however one special case in which the argument still applies: when $\alpha=1$ and $u=\pm 1$. The goal of the two following Propositions is to give sufficient conditions that imply that $\rho_{g}(G_{H})$ contains a non-scalar matrix with $\pm 1$ as an eigenvalue.
@@
special2special Proposition 2.6  Assume that $1\in\varepsilon_{g}(G_{H})$ and that $\rho_{g}(G_{H})$ does not contain an element similar to $\Delta_{-1,1}$. Then $g$ has either real or complex multiplication by some quadratic field $K$. If moreover we can choose $K$ so that it is not contained in $H$, then the following holds:
@@
Proof. –  Let $M\in\operatorname{GL}_{2}(L_{0})$ be a matrix with $\det{M}=-1$ whose image in $\operatorname{PGL}_{2}(L_{0})$ has order $2$. Then $M^{2}$ is scalar, so $M$ is conjugate in $\operatorname{GL}_{2}({\mathbb{C}})$ to $\Delta_{a,b}$ for some $a,b\in{\mathbb{C}}^{\times}$ with a=ba=-b and ab=1ab=-1. Therefore $M$ is similar to $\Delta_{1,-1}$ in $\operatorname{GL}_{2}({\mathbb{C}})$, so $M$ is similar to $\Delta_{1,-1}$ in $\operatorname{GL}_{2}(L_{0})$.
Let $Z\leq\rho_{g}(G_{H})$ be the subgroup of scalar matrices. Consider the surjective determinant map $\delta:\tilde{\rho}_{g}(G_{H})=\rho_{g}(G_{H})/Z\rightarrow\varepsilon_{g}(G_{H})/\det{Z}$, which has cyclic image. Let $x\in\tilde{\rho}_{g}(G_{H})$ be an element of order $2$ such that $\delta(x)=-\det{Z}$: then there is a matrix $M\in\rho_{g}(G_{H})$ above $x$ such that $\det{M}=-1$, so $M$ is similar to $\Delta_{1,-1}$, a contradiction.
@@
Case 1: when $1\notin\det{Z}$.
@@
The image of $\delta$ contains exactly two elements of $2$-torsion: the trivial element $\det{Z}$ and $\det{Z}$. For any $x\in\tilde{\rho}_{g}(G_{H})$ of order two, $\delta(x)\in\mathrm{im}(\delta)$ is a $2$-torsion element which is not $\det{Z}$, so $x\in\ker{\delta}$. Since $\tilde{\rho}_{g}(G_{H})/\ker{\delta}$ contains an element of order two, the subgroup of $\tilde{\rho}_{g}(G_{H})$ generated by its elements of order two has even index: thus, $\tilde{\rho}_{g}(G_{H})$ is not dihedral, and it is not isomorphic to $A_{4}$, $S_{4}$ or $A_{5}$ either. Since $\tilde{\rho}_{g}(G_{H})$ is a subgroup of $\operatorname{PGL}_{2}(L_{0})$, it is cyclic with order $n$ divisible by $4$.
@@
Therefore, the subgroup $\tilde{\rho}_{g}(G_{{\mathbb{Q}}})^{\prime}$ of $\tilde{\rho}_{g}(G_{H})$ is cyclic, so $\tilde{\rho}_{g}(G_{{\mathbb{Q}}})$ is not isomorphic to $A_{4}$, $S_{4}$, or $A_{5}$. Hence $\tilde{\rho}_{g}(G_{{\mathbb{Q}}})$ is dihedral and $g$ has either real or complex multiplication by a quadratic number field $K$. The dihedral group $\tilde{\rho}_{g}(G_{{\mathbb{Q}}})$ contains the cyclic subgroup $\tilde{\rho}_{g}(G_{H})$ of order $n\geq 4$, so $\tilde{\rho}_{g}(G_{H})$ is contained in the unique index two cyclic subgroup of $\tilde{\rho}_{g}(G_{{\mathbb{Q}}})$ (whose inverse image by $\tilde{\rho}_{g}$ is $G_{K}$), so that $K$ is unique and one has $K\subset H$, so we are done.
Case 2: when $1\in\det{Z}$.
@@
In this case, $\rho_{g}(G_{H})$ contains the scalar matrix $iI_{2}$. By the second paragraph of the proof, the group $\ker{\delta}$ does not contain an element of order $2$, so it has odd cardinality, and so does its subgroup $\tilde{\rho}_{g}(G_{{\mathbb{Q}}})^{\prime}$. This is not possible if the group $\tilde{\rho}_{g}(G_{{\mathbb{Q}}})$ is isomorphic to $A_{4}$, $S_{4}$ or $A_{5}$ or if it is dihedral with cardinality divisible by $8$. Hence $\tilde{\rho}_{g}(G_{{\mathbb{Q}}})$ is dihedral with cardinality not divisible by $8$, so in particular, $g$ has either real or complex multiplication by some quadratic number field $K$.
@@
Let us furthermore assume that $K\not\subset H$. Let $C$ denote the cyclic subgroup of $D:=\tilde{\rho}_{g}(G_{{\mathbb{Q}}})$ such that $G_{K}=\tilde{\rho}^{-1}(C)$, and $D_{H}$ denote the subgroup $\tilde{\rho}_{g}(G_{H})$ of $D$. Since $K\not\subset H$, one has $D_{H}\not\subset C$, so $D_{H}$ is dihedral with cyclic subgroup $D_{H}\cap C$ of index $2$ (it is possible, given our assumptions, that $D_{H}\simeq{\mathbb{Z}}/2{\mathbb{Z}}$ and $D_{H}\cap C$ is trivial; this case does not require any special treatment). In particular, the abelianization of $D_{H}$ has exponent $2$, so its cyclic quotient $\mathrm{im}(\delta)$ has at most two elements. Hence either one has $\det{Z}=\varepsilon_{g}(G_{H})$ or one has $\det{Z}=\varepsilon_{g}(G_{H})^{2}$.
@@
Let $x\in D_{H}\backslash C$: then $x$ has order $2$, so $\delta(x)\neq-\det{Z}=\det{Z}$, so the image of $\delta$ has two elements, hence $1\in\det{Z}=\varepsilon_{g}(G_{H})^{2}$, i.e. $i\in\varepsilon_{g}(G_{H})$. Furthermore, $\ker{\delta}$ is a subgroup of index $2$ of $D_{H}$ which has odd cardinality, so the cardinality of $\tilde{\rho}_{g}(G_{H})$ is not divisible by $4$. $\Box$
@@
    $q=3$ divides $H:{\mathbb{Q}}]$, $\tilde{\rho}_{g}(G_{{\mathbb{Q}}})$ is isomorphic to $A_{4}$, $\tilde{\rho}_{g}(G_{H})$ is isomorphic to ${\mathbb{F}}_{2}^{\oplus 2}$.
@@
As in the proof of Proposition 2, the determinant map $\delta:\rho_{g}(G_{H})\rightarrow\varepsilon_{g}(G_{H})/\det{Z}$ is surjective and its kernel contains $Z$ and $\rho_{g}(G_{{\mathbb{Q}}})^{\prime}$. Hence $\delta$ factors through the quotient $\rho_{g}(G_{H})/Z\rho_{g}(G_{{\mathbb{Q}}})^{\prime}\simeq\tilde{\rho}_{g}(G_{H})/\tilde{\rho}_{g}(G_{{\mathbb{Q}}})^{\prime}$: in other words, $\varepsilon_{g}(G_{H})/\det{Z}$ identifies as a cyclic quotient of $\tilde{\rho}_{g}(G_{H})/\tilde{\rho}_{g}(G_{{\mathbb{Q}}})^{\prime}$.
Case 1: $\tilde{\rho}_{g}(G_{{\mathbb{Q}}}):\tilde{\rho}_{g}(G_{{\mathbb{Q}}})^{\prime}]$ is prime to $q$.
In this case, the cardinality of $\tilde{\rho}_{g}(G_{{\mathbb{Q}}})^{\prime}$ is divisible by $q$, so $\tilde{\rho}_{g}(G_{{\mathbb{Q}}})^{\prime}$ contains an element of order $q$. Now, $\tilde{\rho}_{g}(G_{{\mathbb{Q}}})^{\prime}$ is the image in $\operatorname{PGL}_{2}(L_{0})$ of $\rho_{g}(G_{{\mathbb{Q}}})^{\prime}\subset\rho_{g}(G_{H})\cap\operatorname{SL}_{2}(L_{0})$ (the inclusion holds because $H/{\mathbb{Q}}$ is abelian). Hence the image of $\rho_{g}(G_{H})\cap\operatorname{SL}_{2}(L_{0})$ in $\operatorname{PGL}_{2}(L_{0})$ contains an element of order $q$.
@@
Case 2: $\tilde{\rho}_{g}(G_{{\mathbb{Q}}}):\tilde{\rho}_{g}(G_{{\mathbb{Q}}})^{\prime}]$ is divisible by $q$.
If $\tilde{\rho}_{g}(G_{{\mathbb{Q}}})$ is isomorphic to $S_{4},A_{5}$ or is dihedral, then $\tilde{\rho}_{g}(G_{{\mathbb{Q}}})/\tilde{\rho}_{g}(G_{{\mathbb{Q}}})^{\prime}$ is a $2$-group, so its cardinality is not divisible by $q$. Therefore $\tilde{\rho}_{g}(G_{{\mathbb{Q}}})$ is isomorphic to $A_{4}$, so $\tilde{\rho}_{g}(G_{{\mathbb{Q}}})/\tilde{\rho}_{g}(G_{{\mathbb{Q}}})^{\prime}$ is isomorphic to ${\mathbb{Z}}/3{\mathbb{Z}}$, hence one has $q=3$.
@@
Let $M\in\rho_{g}(G_{H})$ be such that $\det{M}$ is a primitive third root of unity. If $M^{2}\in Z$, $M^{4}=\det(M^{2})I_{2}\in Z$ is an element of order $3$, a contradiction: hence the projective image of $M$ has order at least $3$. The elements of $A_{4}$ have order $1$, 22 or $3$, so the projective image of $M$ has order $3$, i.e. $M^{3}\in\operatorname{SL}_{2}(L_{0})$ is scalar, thus $M^{3}\in\{\pm I_{2}\}$. Replacing $M$ with $M^{2}$ if needed, we may assume that $M^{3}=I_{2}$. Let $u,v\in L_{0}$ be the two eigenvalues of $M$, then one has $u^{3}=v^{3}=1$ and $u\notin\{v,v^{-1}\}$ (since $M$ is not scalar and $\det{M}$ is a primitive third root of unity), so exactly one of $u$ and $v$ is equal to $1$, and we are done.
@@
By Case 2 (a), we may assume that $Z$ (and therefore $\det{Z}$) contains an element of order $3$, so it is enough to show that there exists $M\in\rho_{g}(G_{H})\cap\operatorname{SL}_{2}(L_{0})$ with projective image of order $3$. The cyclic group $\varepsilon_{g}(G_{H})$ contains no element of order $9$, so the quotient $\varepsilon_{g}(G_{H})/\det{Z}$ has cardinality coprime to $3$. The abelianization of $\tilde{\rho}_{g}(G_{H})\simeq A_{4}$ is isomorphic to ${\mathbb{Z}}/3{\mathbb{Z}}$, so $\delta$ is trivial and surjective, hence $\varepsilon_{g}(G_{H})=\det{Z}$. This implies directly that $\rho_{g}(G_{H})\cap\operatorname{SL}_{2}(L_{0})\rightarrow\tilde{\rho}_{g}(G_{H})$ is surjective. We are done, because $\tilde{\rho}_{g}(G_{H})\simeq A_{4}$ has an element of order $3$.
@@
sufficient-for-euler-adaptedspecial Corollary 2.8  Let $\mathfrak{p}$ be a good prime for $(f,g)$. If $\rho_{g}(G_{H})$ contains a non-scalar matrix $M$ such that $1$ or 1-1 is an eigenvalue of $M$, then $T_{f,g,\mathfrak{p}}$ satisfies (wE); if furthermore the order of $M$ is prime to $p$, then $T_{f,g,\mathfrak{p}}$ satisfies (sE). Hence $T_{f,g,\mathfrak{p}}$ has Euler-adapted Galois image for all but finitely many primes $\mathfrak{p}$. In particular, all of the above applies if any of the following conditions is satisfied:
@@
    sc2special $1\in\varepsilon_{g}(G_{H})$ and $\tilde{\rho}_{g}(G_{{\mathbb{Q}}})$ is not dihedral.
@@
    sc3special $1\in\varepsilon_{g}(G_{H})$, $\tilde{\rho}_{g}(G_{H})$ is a dihedral group of order $4n$ with $n\geq 1$.
@@
    sc4special $1\in\varepsilon_{g}(G_{H})$, $\tilde{\rho}_{g}(G_{{\mathbb{Q}}})$ is a dihedral group of order $8n$ with $n\geq 1$, and $g$ has either real or complex multiplication by some quadratic field $K\not\subset H$.
@@
    sc5special $\varepsilon_{g}(G_{H})$ contains 1-1 but not $i$, and, for some quadratic number field $K$ by which $g$ has either real or complex multiplication, one has $K\not\subset H$.
@@
Proof. –  By Proposition 2, we only need to discuss conditions (wE) and (sE). Let $\mathfrak{p}$ be a good prime with residue characteristic $p$ and $M\in\rho_{g}(G_{H})$ be a non-scalar matrix admitting $s\in\{\pm 1\}$ as an eigenvalue. By Proposition 2, there exists $\sigma\in G_{H(\mu_{p^{\infty}})}$ such that $\rho_{g}(\sigma)=M$ and $\sigma$ acts on $T_{f,\mathfrak{p}}$ by $sU$. Then the kernel of the action of $\sigma-1$ on $T_{f,g,\mathfrak{p}}\otimes(L_{0})_{\mathfrak{p}}$ is a line and $T_{f,g,\mathfrak{p}}$ satisfies (wE). If the order of $M$ is prime to $p$, then $\mathcal{O}_{(L_{0})_{\mathfrak{p}}}^{\oplus 2}/(M-sI_{2})\mathcal{O}_{(L_{0})_{\mathfrak{p}}}^{\oplus 2}$ is a free $\mathcal{O}_{(L_{0})_{\mathfrak{p}}}$-module of rank one, so $T_{f,g,\mathfrak{p}}/(\sigma-1)T_{f,g,\mathfrak{p}}$ is a free $\mathcal{O}_{(L_{0})_{\mathfrak{p}}}$-module of rank one, thus $T_{f,g,\mathfrak{p}}$ satisfies (sE).
All that remains to do is show that in conditions (i)-(vi), there exists a non-scalar matrix $M\in\rho_{g}(G_{H})$ which has $1$ or 1-1 as an eigenvalue.